# A Novel Algorithm for Digital Lithological Mapping: Case Studies in Sri Lanka’s Mineral Exploration

### v1

@@ 4.6.1 Magnetic Seperation Method
Ilmenite is a mineral that possesses some magnetic susceptibility. Hence, it is possible to identify the availability of Ilmenite in collected specimens incorporating magnetic methods. In this study, FRANTZ Magnetic Separator Model L-1 was used to separate Ilmenite. Prior to magnetic separation, the samples were prepared according to the steps described below.
@@
    The standard particle size for magnetic separation is 75µm - 125µm. Through the sieving process, the samples with the preferred grain size were separated.
@@
    As the magnetic separator requires magnetite removal prior to the samples being entered into the separator, a hand magnet was used to separate the magnetite mineral from the obtained soil samples.
@@ 4.6.2 X-Ray Diffraction Method
X-Rays are generated as a result of the deceleration of electrically charged particles which contain a sufficient amount of energy. The XRD test which incorporates these is considered to be a non destructive technique that unveil crystallographic structure, physical properties and chemical composition of a mineral. The availability of a certain mineral is detected through the analysis of diffracted rays, which are generated by directing collimated X-Rays towards the sample. Before carrying out the XRD test a pre-processing procedure which is unique to each mineral was followed. The main common steps followed for sample preparation is listed below.
@@
    Representative samples were separated form each main sample from the sites.
@@
    Powdered sample was filtered for 63 microns through the sieve shake.
@@
For Limestone the following analysis steps were conducted before sample preparation.
@@
    The texture ,color and grain size of the samples were manually inspected and logged.
@@
    A total of twenty samples were selected (five samples from each site) for XRD. The selection was done accounting locations with high and low abundance values.
@@ 5.1 Identification of the Number of Endmembers
As described in section 4.1, the image data of a particular terrain underwent iterative clustering by varying the number of clusters from one to ten. To determine the number of end-members using the elbow method, the within-cluster sum of squares (WCSS) was calculated for each iteration, and a plot of WCSS vs. Number of Clusters was generated. Typically, the optimal number of clusters for the data is considered to be the number corresponding to the elbow point on the plot. The plots for the four terrains are illustrated in Fig. 4. After identifying the elbow points, the number of dominant endmembers present in each location was determined and tabulated in Table 4.
Figure 4: Variation of the WCSS with the number of Clusters for each Region: (a) Jaffna (b) Pulmoddai (c) Mannar (d) Giant’s Tank
@@
From Fig. 4, it can be seen that the WCSS value drops significantly for Mannar, Pulmoddai and Jaffna when the number of clusters were set to two. In comparison, the WCSS of the Giant’s Tank does not drop as significantly. When considering the WCSS plots in general, a significant drop in the WCSS value demarcate the separation of a large cluster with a low variance. In this context the above observation for Mannar, Pulmoddai and Jaffna is a result of the separation of highly abundant water pixels present in the HSIs of those regions. On the contrary, in the Giant’s Tank’s HSI, the absence of such a large cluster of low variance could be attributed to the insignificant difference in the WCSS after the formation of two clusters. Therefore, one can get a sense of the presence of an abundant endmember with low variance in a given region by observing the WCSS plots.
@@ 5.2 Identification and Isolation of Macroscopical Components
With insights gained from the elbow method regarding the number of endmembers, VCA was employed by the authors to extract the dominant end-members from each hyperspectral image (HSI). The extracted endmembers are visually depicted in Fig. 5. In order to identify and isolate pixels corresponding to each candidate endmember, the authors utilized the similarity criteria mentioned in section 4.1. It is evident that the pre-classification stage successfully captured the land profile, effectively segmenting areas corresponding to different macroscopical components. Consequently, the identified end-members were labeled with their respective macroscopical elements, as summarized in table 4.
From the generated maps, it is evident that the algorithm has been able to identify the regions corresponding to each endmember, allowing the isolation of soil pixels, which are used to perform further analysis for single target mineral detection. While the proposed algorithm for the pre-classification block performed admirably, it is imperative to explore the adjustability and scalability of the proposed structure as some applications would require more refined extraction of end members and identification of the land cover of the particular endmember.
Figure 5: The Macroscopic Endmember signatures extracted for the four locations considered: (a) Jaffna (b) Pulmoddai (c) Mannar (d) Giant’s Tank
Figure 6: The Mineral and Impurity representative signatures and the Laboratory Mineral signature for each location : (a) Jaffna (b) Pulmoddai (c) Mannar (d) Giant’s Tank
@@ 5.3 Correlation Analysis on Soil Pixels
Once the pixels belonging to the soil subclass were isolated using the aforementioned methodology, their correlation with laboratory mineral signatures obtained from the USGS Spectral Library was calculated. In accordance with section 3.4, the extracted endmembers from the spectral library underwent several pre-processing steps to generate the signatures corresponding to Hyperion wavelengths. These pre-processed signatures provide a valuable reference for comparison and analysis in the subsequent stages of the study.
To determine the representative classes for each study, the first step involved extracting two signatures from the soil manifold. Subsequently, classification was carried out using the algorithm outlined in section 4.2. In this classification process, the pixels were distinguished based on their correlation with the laboratory mineral signature. The range of correlation values for each study, along with the upper and lower bounds necessary for identifying the representative classes, are summarized in table 5. This table provides an overview of the correlation thresholds used to assign pixels to the mineral or impurity representative classes.
@@
Based on the range of correlation values, it is evident that the laboratory signature of Montmorillonite has a positive correlation with pixel signatures belonging to a specific soil subclass in the area. This correlation is supported by the positive correlation observed between the laboratory signature and the two representative signatures extracted from the soil subclass. However, in the Jaffna terrain, the soil pixels are more likely to exhibit a negative correlation with the laboratory signature of Limestone. Furthermore, the correlation range for soil pixels in Pulmoddai and Mannar, with respect to the Ilmenite reference signature, is quite similar. These findings align with the mineralogical background discussed in the corresponding section.
@@
To develop the parameter, relative availability, the eigendirection which enhances the separation between the two subclasses was computed using Fisher’s Discriminant Analysis for each area of focus. Fig. 7 show the presence of the two subclasses in the reduced space.
Figure 7: Mineral and Impurity separaton in the reduced domain: (a) Jaffna (b) Mannar (c) Giant’s Tank (d) Pulmoddai
From the above figures, it can be seen that FDA has been able to find an eigen-direction such that there is a good amount of separability between the two subclasses.
It can be observed that the Montmorillonite pixels, after being projected to the eigen-direction, show the largest relative separation (0.1) between the mean of the two subclasses. However, the other three regions’ signatures do not show a separability (0.02 - Limestone, 0.01 - Mannar) in comparison to the Montmorillonite pixels. Moreover, it can be seen that the two clusters representing mineral representative and impurity representative pixels corresponding to Limestone and Pulmoddai show a reasonable separation in relation to the variance of the clusters. On the contrary, the Fig. 7 corresponding to Mannar depicts a slight overlap between the two representative clusters. This might be due to the fact that the site contains a low concentration of the mineral. Having less mineral presence in a given pixel increases the presence of the impurities pushing the mineral subclass closer to the impurity subclass. This claim is proven by the XRD validation of the samples collected from the Mannar region. The samples collected from the region show less mineral presence which has caused the two subclasses to be close and slightly overlapping.
@@ 5.4.1 Mannar - Ilmenite
Table 6: XRD Test Results, Relative Availability and Abundance Estimation after NNLS for the Sites at Mannar
@@ 5.4.2 Pulmoddai
Site Magnetic Separation Test Results Relative Availability Abundance after NNLS Estimate
@@ 5.4.3 Jaffna
The provided Fig.11 illustrates the soil characteristics and surface texture of four survey sites. To begin with, site 1 (Fig.10(a)) was characterized by sparse vegetation and had dry and compact soil, without any visible signs of Limestone in the vicinity. Moving on to site 3 (Fig.10(b)), there was no apparent presence of surface Limestone, and the reddish-brown soil color differed from the other three sites. However, it’s worth noting that Site-3 was situated in a populated area and had experienced ongoing human activity, which may have led to abrupt changes in soil composition, differing from what would be expected based on historical geological conditions.In contrast, site 2 (Fig.10(c)) featured moist, dark brown soil and exhibited seashell fragments scattered on the surface. This area was identified as a drained lagoon, and it was large enough to be discerned through satellite imagery. The widespread presence of seashell remnants throughout the lagoon area suggested their deposition due to tides, serving as evidence for the existence of Limestone on the surface.
Figure 10: In-situ soil charateristic observations for (a) site-1, (b) site-3, (c) site-2, and (d) site-4
@@
To buttress the observations from these sites, the amount of Limestone available according to the XRD test is provided in Table 8. As tabulated in Table 8, the proposed method had made an accurate prediction of the Limestone presence in site 2 and site 4 with higher $\alpha$ values while returning low $\alpha$ values for site 1 and site 3 that implies less chance of Limestone presence, which is in agreement with the observations made by the validation field survey via XRD tests. Additionally, the correlation coefficient between the XRD values and the relative availability values resulted in a value of 0.9853. This further confirms the fact that algorithm has been able to accurately predict the availibility in these sites. Furthermore, the relative availability and the alpha values calculated for the four sites show a good correspondence with the XRD test results. Based on the results it is evident that the algorithm performs admirably across sites with different soil characteristics to find feasible survey sites for Limestone availability.
@@ 5.4.4 Giant’s Tank
The availability of Montmorillonite in the Giant’s Tank and Murunkan regions in Sri Lanka has been confirmed by earlier research (Wanasinghe and Adikary, (2017); Herath, (1973)). Montmorillonite, being a clay mineral, is abundant in arid regions, often found in tanks constructed for cultivation purposes. Geomorphologically, Murunkan primarily features flat terrain with numerous shallow water bodies nearby. These water bodies periodically dry up during extended droughts but are replenished mainly by rainwater. The clay-rich composition of the soil contributes to a high surface water retention capacity due to its low infiltration rate. Montmorillonite tends to accumulate at the bottom of these tanks through sedimentation and is also commonly found in the regions surrounding the tanks and the fields irrigated using water from these reservoirs.
This study specifically focuses on the region containing the Giant’s Tank, which is a prominent irrigation tank located in Northern Sri Lanka. This region is characterized by abundant wetlands and paddy fields, likely containing traces of Montmorillonite.
Table 9: XRD Test Results, Relative Availability and Abundance Estimation after NNLS for the Sites at Giant’s Tank

### v2

@@ 4.6.1 Magnetic Seperation Method
Ilmenite is a mineral that possesses some magnetic susceptibility. Hence, it is possible to identify the availability of Ilmenite in collected specimens incorporating magnetic methods. This study used FRANTZ Magnetic Separator Model L-1 to separate Ilmenite. Prior to magnetic separation, the samples were prepared according to the steps described below.
@@
    The standard particle size for magnetic separation is 75µm - 125µm. The samples with the preferred grain size were separated through the sieving process.
@@
    As the magnetic separator requires magnetite removal before the samples are entered into the separator, a hand-magnet was used to separate the magnetite mineral from the obtained soil samples.
@@ 4.6.2 X-Ray Diffraction Method
X-rays are generated due to the deceleration of electrically charged particles, which contain a sufficient amount of energy. The XRD test, which incorporates these, is considered to be a non-destructive technique that unveils the crystallographic structure, physical properties and chemical composition of a mineral. The availability of a particular mineral is detected through the analysis of diffracted rays, which are generated by directing collimated X-rays towards the sample. Before carrying out the XRD test, a pre-processing procedure unique to each mineral was followed. The main common steps followed for sample preparation are listed below.
@@
    Representative samples were separated from each main sample from the sites.
@@
    The powdered sample was filtered for 63 microns through the sieve shake.
@@
For Limestone, the following analysis steps were conducted before sample preparation.
@@
    The samples’ texture, colour and grain size were manually inspected and logged.
@@
    A total of twenty samples were selected (five samples from each site) for XRD. The selection was done by accounting for locations with high and low abundance values.
@@ 5.1 Identification of the Number of Endmembers
As described in section 4.1, the image data of a particular terrain underwent iterative clustering by varying the number of clusters from one to ten. To determine the number of endmembers using the elbow method, the within-cluster sum of squares (WCSS) was calculated for each iteration, and a plot of WCSS vs. Number of Clusters was generated. Typically, the optimal number of clusters for the data is the number corresponding to the elbow point on the plot. The plots for the four terrains are illustrated in Fig. 4. After identifying the elbow points, the number of dominant endmembers in each location was determined and tabulated in Table 4.
(a) Jaffna
(b) Pulmoddai
(c) Mannar
(d) Giant’s Tank
Figure 4: Variation of the WCSS with the number of Clusters for each Region
@@
From Fig. 4, it can be seen that the WCSS value dropped significantly for Mannar, Pulmoddai and Jaffna when the number of clusters was set to two. In comparison, the WCSS of the Giant’s Tank does not drop as significantly. When considering the WCSS plots in general, a significant drop in the WCSS value demarcates the separation of a large cluster with a low variance. In this context, the above observation for Mannar, Pulmoddai and Jaffna results from the separation of highly abundant water pixels in the HSIs of those regions. On the contrary, in the Giant’s Tank’s HSI, the absence of such a large cluster of low variance could be attributed to the insignificant difference in the WCSS after forming two clusters. Therefore, one can get a sense of the presence of an abundant endmember with low variance in a given region by observing the WCSS plots.
@@ 5.2 Identification and Isolation of Macroscopical Components
With insights gained from the elbow method regarding the number of endmembers, VCA was employed by the authors to extract the dominant endmembers from each hyperspectral image (HSI). The extracted endmembers are visually depicted in Fig. 5. In order to identify and isolate pixels corresponding to each candidate endmember, the authors utilized the similarity criteria mentioned in section 4.1. It is evident that the pre-classification stage successfully captured the land profile, effectively segmenting areas corresponding to different macroscopical components. Consequently, the identified endmembers were labelled with their respective macroscopical elements, as summarized in Table 4.
The generated maps show that the algorithm has identified the regions corresponding to each endmember, allowing the isolation of soil pixels, which are used to perform further analysis for single-target mineral detection. While the proposed algorithm for the pre-classification block performed admirably, it is imperative to explore the adjustability and scalability of the proposed structure as some applications would require more refined extraction of end members and identification of the land cover of the particular endmember.
(a) Jaffna
(b) Pulmoddai
(c) Mannar
(d) Giant’s Tank
Figure 5: The Macroscopic Endmember signatures extracted for the four locations considered
(a) Jaffna
(b) Pulmoddai
(c) Mannar
(d) Giant’s Tank
Figure 6: The Mineral and Impurity representative signatures and the Laboratory Mineral signature for each location
@@ 5.3 Correlation Analysis on Soil Pixels
Once the pixels belonging to the soil subclass were isolated using the methodology mentioned earlier, their correlation with laboratory mineral signatures obtained from the USGS Spectral Library was calculated. Per section 3.4, the extracted endmembers from the spectral library underwent several pre-processing steps to generate the signatures corresponding to Hyperion wavelengths. These pre-processed signatures provide a valuable reference for comparison and analysis in the subsequent stages of the study.
To determine the representative classes for each study, the first step involved extracting two signatures from the soil manifold. Subsequently, classification was performed using the algorithm outlined in section 4.2. In this classification process, the pixels were distinguished based on their correlation with the laboratory mineral signature. The range of correlation values for each study and the upper and lower bounds necessary for identifying the representative classes are summarised in Table 5. This table overviews the correlation thresholds for assigning pixels to the mineral or impurity representative classes.
@@
Based on the range of correlation values, it is evident that the laboratory signature of Montmorillonite has a positive correlation with pixel signatures belonging to a specific soil subclass in the area. This correlation is supported by the positive correlation between the laboratory signature and the two representative signatures extracted from the soil subclass. However, in the Jaffna terrain, the soil pixels are more likely to correlate negatively with Limestone’s laboratory signature. Furthermore, the correlation range for soil pixels in Pulmoddai and Mannar, with respect to the Ilmenite reference signature, is quite similar. These findings align with the mineralogical background discussed in the corresponding section.
@@
To develop the parameter relative availability, the eigendirection, which enhances the separation between the two subclasses, was computed using Fisher’s Discriminant Analysis for each area of focus. Fig. 7 show the presence of the two subclasses in the reduced space.
(a) Jaffna
(b) Mannar
(c) Giant’s Tank
(d) Pulmoddai
Figure 7: Mineral and Impurity separaton in the reduced domain
From the above figures, it can be seen that FDA has found an eigen-direction such that there is a good amount of separability between the two subclasses.
It can be observed that the Montmorillonite pixels after being projected to the eigen-direction, show the largest relative separation (0.1) between the mean of the two subclasses. However, the signatures of the other three regions do not show separability (0.02 - Limestone, 0.01 - Mannar) compared to the Montmorillonite pixels. Moreover, it can be seen that the two clusters representing mineral representative and impurity representative pixels corresponding to Limestone and Pulmoddai show a reasonable separation in relation to the variance of the clusters. On the contrary, Fig. 7 corresponding to Mannar depicts a slight overlap between the two representative clusters. This might be because the site contains a low concentration of the mineral. Less mineral presence in a given pixel increases the presence of the impurities, pushing the mineral subclass closer to the impurity subclass. This claim is proven by the XRD validation of the samples collected from the Mannar region. The samples collected from the region show less mineral presence, which has caused the two subclasses to be close and slightly overlapping.
@@ 5.4.1 Mannar - Ilmenite
Table 6: XRD Test Results, Relative Availability and Abundance Estimation after Non-negative Least Squares (NNLS) for the Sites at Mannar
@@ 5.4.2 Pulmoddai
Site Magnetic Separation Relative Availability Abundance after
Test Results NNLS Estimate
@@ 5.4.3 Jaffna
The provided Fig.11 illustrates the soil characteristics and surface texture of four survey sites. To begin with, site 1 (Fig.10(a)) was characterised by sparse vegetation and had dry and compact soil, without any visible signs of Limestone in the vicinity. Moving on to site 3 (Fig.10(b)), there was no apparent presence of surface Limestone, and the reddish-brown soil colour differed from the other three sites. However, it’s worth noting that Site-3 was situated in a populated area and had experienced ongoing human activity, which may have led to abrupt changes in soil composition, differing from what would be expected based on historical geological conditions.In contrast, site 2 (Fig.10(c)) featured moist, dark brown soil and exhibited seashell fragments scattered on the surface. This area was identified as a drained lagoon, and it was large enough to be discerned through satellite imagery. The widespread presence of seashell remnants throughout the lagoon area suggested their deposition due to tides, serving as evidence for the existence of Limestone on the surface.
(a) site-1
(b) site-3
(c) site-2
(d) site-4
Figure 10: In-situ soil characteristic observations
@@
To buttress the observations from these sites, the amount of Limestone available according to the XRD test is provided in Table 8. As tabulated in Table 8, the proposed method had made an accurate prediction of the Limestone presence in site 2 and site 4 with higher $\alpha$ values while returning low $\alpha$ values for site 1 and site 3 that implies less chance of Limestone presence, which is in agreement with the observations made by the validation field survey via XRD tests. Additionally, the correlation coefficient between the XRD values and the relative availability values resulted in a value of 0.9853. This further confirms the fact that algorithm has been able to accurately predict the availability in these sites. Furthermore, the relative availability and the alpha values calculated for the four sites show a good correspondence with the XRD test results. Based on the results it is evident that the algorithm performs admirably across sites with different soil characteristics to find feasible survey sites for Limestone availability.
@@ 5.4.4 Giant’s Tank
The availability of Montmorillonite in the Giant’s Tank and Murunkan regions in Sri Lanka has been confirmed by earlier research (Wanasinghe and Adikary 2017); (Herath 1973). Montmorillonite, being a clay mineral, is abundant in arid regions, often found in tanks constructed for cultivation purposes. Geomorphologically, Murunkan primarily features flat terrain with numerous shallow water bodies nearby. These water bodies periodically dry up during extended droughts but are replenished mainly by rainwater. The clay-rich composition of the soil contributes to a high surface water retention capacity due to its low infiltration rate. Montmorillonite tends to accumulate at the bottom of these tanks through sedimentation and is also commonly found in the regions surrounding the tanks and the fields irrigated using water from these reservoirs.
This study specifically focuses on the region containing the Giant’s Tank, which is a prominent irrigation tank located in Northern Sri Lanka. This region is characterised by abundant wetlands and paddy fields, likely containing traces of Montmorillonite.
Table 9: XRD Test Results, Relative Availability, and Abundance Estimation after NNLS for the Sites at Giant’s Tank